OpenStudy (anonymous):

what is 2 radical 5 (in parenthasees) squared equal to and how do you solve it?

OpenStudy (anonymous):

\[(2\sqrt{5})^{2} = 2 \times 2 \times \sqrt{5} \times \sqrt{5}\] Try this!